OpenStudy (anonymous):

Find the linearization of the function f(x,y)=sqrt(34−x^2−2y^2) at the point (-1, 2).

OpenStudy (anonymous):

I got Z = x/10 + 4y/5 + 35/10 as my answer, where did I go wrong?

OpenStudy (anonymous):

tried again, got z = x/10 - 4y/5 + 67/10... Still wrong... Any ideas Robtobey?

OpenStudy (anonymous):

you guys here?

OpenStudy (anonymous):

first f(-1,2) = 5 now f(x,y) = f(-1,2) + fx(-1,2)(x-(-1)) + fy(-1,2)(y-2)

OpenStudy (anonymous):

so fx = (1/2)[(34−x^2−2y^2)^(-1/2)] * (-2x) = -x * (34−x^2−2y^2)^(-1/2) so fx(-1,2) = 0.2

OpenStudy (anonymous):

how did I get my fx(-1,2) to be .1?

OpenStudy (anonymous):

you ask me ? :P

OpenStudy (anonymous):

that answer you gave didn't work, also :(

OpenStudy (anonymous):

yes i have an error

OpenStudy (anonymous):

fy = (1/2)[(34−x^2−2y^2)^(-1/2)] *(-4y) = -2y * (34−x^2−2y^2)^(-1/2) fy(-1,2) = -4 * 0.2 = -0.8 so finally f(x,y) = 5 + 0.2(x+1)-0.8(y-2) f(x,y) = 0.2x -0.8y + 6.8

OpenStudy (anonymous):

fixed here .. i deleted the wrong

OpenStudy (anonymous):

hmm Thanks, I'll compare your work to mine to see that arithmatic I fluttered up. thanks again man.

OpenStudy (anonymous):

the last one is good right ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok :)